dx/dt=2x+2y, dy/dt=x+3y



1
Expert's answer
2021-02-23T11:20:41-0500

Want to solve the following system:,

"\frac{dx}{dt}=2x+2y\hspace{10pt}(1)\\
 \frac{dy}{dt}=x+3y\hspace{15pt}(2)"

i.e., want to find "x(t)" and "y(t)" which satisfy the above two equations.


Differentiating equation (1) with respect to "t" yields,

"\frac{d^2 x}{dt^2}=2\frac{dx}{dt}+2\frac{dy}{dt}"


Substituting "\frac{dy}{dt}" from equation (2), we obtain,

"\frac{d^2x}{dt^2}=2\frac{dx}{dt}+(2x+6y)"


Substituting "y" from equation (1) gives,

"\frac{d^2x}{dt^2}=2\frac{dx}{dt}+2x+3\left(\frac{dx}{dt}-2x\right)\\
\Rightarrow \frac{d^2x}{dt^2}=5\frac{dx}{dt}-4x\\
\Rightarrow \frac{d^2x}{dt^2}-5\frac{dx}{dt}+4x = 0\hspace{20pt}(3)"


Clearly equation (3) has a solution of the form,

"x(t)=Ae^{r_1 t}+Be^{r_2 t}" , where "A, B" are two fixed constants and "r_1" and "r_2" are the two roots of the equation:

"r^2 - 5r+4=0\\
 \Rightarrow (r-1)(r-4)=0\\
 \Rightarrow r=1\text{ or }r=4"


Thus we can take "r_1=1" and "r_2=4"

Hence, "x(t)=Ae^t + Be^{4t}"


Using this value of "x(t)" , we get "y(t)" (from equation (1)) as follows,

"y(t)=\frac{1}{2}\left(\frac{dx}{dt}\right)-x\\
 \Rightarrow y(t)=-\frac{A}{2}e^t + B e^{4t}"


So the solution of the system is given by,

"x(t)=Ae^t + Be^{4t}\\
 y(t)=-\frac{A}{2}e^t + B e^{4t}"


The values of "A" and "B" has to be determined from the initial values, if provided (for example, those can be obtained if "x(0)" and "y(0)" are known.)